OpenStudy (4meisu):

10x+ 6y + z = 7 5x - 9y - 2z = 3 15x -12y + 2z = -5 Solve x, y, z by elimination.

OpenStudy (anonymous):

So you woukd solve using 2 equations and then you would use the answer to to solve the third equation

OpenStudy (anonymous):

So you would multiply the second equation by 2 so that the first terms of both equations will equal the same. so here it would be 10x. 2(5x-9y-2z=3) =10x-18y-4z=6 and then you would do elimination Subtract the two equations so you would eliminate one variable 10x-10x=0 6y-(-18y)=24y z-(-4z)=5z 7-6=1 So the new equation would be 24y+5z=1 then you would adsubtract the next two equations to eliminate the x also so you must multipy the equation by 3 for the first variables to equal: 3(5x-9y-2z=3)=15x-27y-6z=9

OpenStudy (anonymous):

Now you must eliminate the last two equations: 15x-15x=0x -27y-(-12y)=-15y -6z-2z=-8z 9-(-5)=14 So we are left with these two equations: -15y-8z=14 and 24y+5z=1 So I guess its easier to eliminate the z. So we must multiply both quation to get the same variable; 5(-15y-8z=14)= -75y-40z=70 and 8(24y+5z=1)=192y+40z=8

OpenStudy (anonymous):

Now we will do elimination: 192y+75y)=267y 40z+(-40z)=0z 8+70=78 267y=78 y=78/267 y=26/89 I must have done wrong calculations becuase I am getting bad numbers

OpenStudy (4meisu):

I used my graphic calculator and y is supposed to be 2/3 :/

OpenStudy (anonymous):

Basically you eliminate a variable in the first two equations. Then you eliminate the same variable in the second two equations. With these two new equations you eliminate a 2nd variable and then you solve fory and then.. I am doing it in my head now so I may have made a mistake I will do it by hand and I will get back to you in a few minutes

OpenStudy (4meisu):

Oka, thank you so much!

OpenStudy (anonymous):

I got the answer. I will begin from the beginning

OpenStudy (4meisu):

Oka, thank you so much!

OpenStudy (4meisu):

alright then

OpenStudy (anonymous):

So we will eliminate the z in all 3 equation. You must multiply the first equation by 2: 2(10x+6y+z=7)= 20x+12y+2z=14 You must add the first two equations to eliminate z: 20x+5x=25x 12y+(-9y)=3y 2z+(-2z)=0z 14+3=17 -->25x+3y=17 Now Eliminate the last two equations: 5x+15x=20x -9y+(-12y)=-21y -2z+2z=0z 3+(-5)=-2 --> 20x-21y=-2 Now you eliminate the y in the two new equations: First you must multiply one equation by 7 for the ys to be eliminated: 7(25x+3y=17)= 175x+21y=119 175x+20x=195x 21y+(-21y)=0y 119+(-2)=117 ---> 195x=117 117/195=x x=3/5

OpenStudy (anonymous):

Now you must plug in the x-value into one of the 2 previous equation to find out what y equals: 25(3/5)+3y=17 3y=17-15 3y=2 y=2/3 Then you plug in the x and y value into one of the three original equations: 10(3/5)+6(2/3)+z=7 6+4+z=7 z=-3 So x=3/5 y=2/3 and z=-3 It took way longer than I had expected it to take but here is the answer. I am a terribly slow typer when it comes to math equations

OpenStudy (4meisu):

Oka, thank you so much!

OpenStudy (4meisu):

Thank you so much!

OpenStudy (4meisu):

Oka, thank you so much!

OpenStudy (4meisu):

this really helped